a spring scale is used to exert a net force of 2.7 N on a cart.if the cart’s mass is 0.64 kg,what is the cart’s acceleration
1
Expert's answer
2020-02-18T10:12:46-0500

Solution. Using Newton’s second law, we write


"F=ma"

where F=2.7N is force; m=0.64kg is mass; a is acceleration. Therefore


"a=\frac{F}{m}=\frac{2.7}{0.64}\approx 4.22\frac{m}{s^2}"

Answer.

"a=4.22\frac{m}{s^2}"